# ERDŐS–PÓSA PROPERTY OF CYCLES THAT ARE FAR APART

### v2

@@ 3 The Proof
The only remaining possibility is that $v^{\prime}$ is an internal vertex of $P_{x^{\prime}}$ for some $x^{\prime}\in I$. Since every internal vertex of $P_{x^{\prime}}$ has degree $2$ in $G^{\prime}$, this implies that $P_{x^{\prime}}\subseteq D^{\prime}$. Let $w^{\prime}$ be the endpoint of $P_{x^{\prime}}$ in $C_{1}$. Since $x^{\prime}\in I\subseteq V(H)$, $x^{\prime}$ is interesting. Therefore $P_{x^{\prime}}\subseteq P_{1,x^{\prime}}\cup P_{2,x^{\prime}}$ contains no vertex in $Y_{1}\cup Y_{2}$.
@@
First suppose that $v^{\prime}\in V(Q^{\prime})$. Let $Q$ be the subpath of $Q^{\prime}$ from $w^{\prime}$ to $v^{\prime}$. Since $Q\subseteq Q^{\prime}\subseteq P_{x^{\prime}}$ contains no vertex in $Y_{1}$ and $V(Q^{\prime})\subseteq B_{G}(V(C_{1}),r)$, we have that $Q\subseteq U_{1}$. Since $U_{1}$ is a $C_{1}$-rooted spanning BFS-unicycle of $G_{1,r}$, $\operatorname{dist}_{G}(w^{\prime},v^{\prime})=\operatorname{dist}_{U_{1}}(w^{\prime},v^{\prime})=\operatorname{len}(Q)$. Observe that $Q$ is a shortest path in $G$ between $V(C_{1})$ and $v^{\prime}$, and thus $\operatorname{len}(Q)\leq\operatorname{len}(P)\leq d$. It follows that $Q\subseteq G[B_{G}(V(C_{1}),d)]$. Since $C_{1}$ is $d$-unicyclic, all paths from $V(C_{1})$ to $v^{\prime}$ in $G[B_{G}(V(C_{1}),d)]$ contain $w^{\prime}$. In particular, $P\subseteq G[B_{G}(V(C_{1}),d)]$ is a path from $v\in V(C_{1})$ to $v^{\prime}$, so $w^{\prime}\in V(P)$. Since $V(P)\cap V(C_{1})=\{v\}$, this implies that $v=w^{\prime}$. However $v=w^{\prime}\in V(P_{x^{\prime}})\subseteq V(D^{\prime})$. This implies that $v$ is a vertex of both $D$ and $D^{\prime}$, which contradicts the fact that $D$ and $D^{\prime}$ are vertex disjoint.
@@
$\geq 2(r+1)-2d+(2d+1)\geq 2r+3\enspace.$
(The second inequality follows from $\operatorname{dist}_{G}(V(C_{1}),q^{\prime})\leq\operatorname{dist}_{G}(V(C_{1}),v^{\prime})+\operatorname{dist}_{G}(v^{\prime},q^{\prime})$ and $\operatorname{dist}_{G}(V(C_{1}),V(C_{2}))\leq\operatorname{dist}_{G}(V(C_{1}),v^{\prime})+\operatorname{dist}_{G}(v^{\prime},V(C_{2}))$.) This contradicts the fact that $\operatorname{len}(P_{x^{\prime}})\leq\operatorname{len}(P_{1,x^{\prime}})+\operatorname{len}(P_{2,x^{\prime}})\leq 2r$. This completes the proof (assuming $|I|\geq\tfrac{1}{2}s(k)$) that outcome a holds.
@@
$X:=\bigcup_{y\in I}\left(V(P_{y})\cap(V(C_{1})\cup V(C_{2}))\right)\enspace,$
be the set of degree-$3$ vertices in $G^{\prime}$. Note that $|X|=2\cdot|I|<s(k)$. We show that $X$ and $Y$ satisfy b.
Consider some vertex $y\in Y$. If $y$ is a $U_{i}$-descendant of some vertex in $Y_{i}$ for some $i\in[2]$, then $y$ satisfies Item x. Therefore, we assume that this is not the case, which implies that $y$ is interesting, so $y\in V(H)$. Since $I$ is a maximal independent set in $H$, the set $I$ is also a dominating set. Therefore, either $y\in I$ or $y$ has a neighbor in $I$. In $G$, this translates into the statement: there exists $y^{\prime}\in I$ such that $\operatorname{dist}_{G}(V(P_{y}),V(P_{y^{\prime}}))\leq d$. (This includes the case $y=y^{\prime}$, which applies when $y\in I$.)
@@
We have just shown that $X$ and $Y$ satisfy ii. Furthermore, $Y=V(G_{1,r})\cap V(G_{2,r})$. Thus, the sets $X$ and $Y$ satisfy the conditions of outcome b. ∎
Proof of Lemma 5.
Let $\mathcal{C}$ be a $2d$-packing of $d$-unicyclic cycles in $G$. If $|\mathcal{C}|\geq k$, then a holds. Thus, let $\mathcal{C}:=\{C_{1},\ldots,C_{p}\}$ and we assume that $p<k$.
For each $i\in[p]$, let $U_{i}$ be a spanning BFS-unicycle of $G[B_{G}(V(C_{i}),r)]$ and let $Y_{i}\subseteq V(G)$ and $X_{i}\subseteq V(G)$ be the sets obtained by applying Lemma 6 to $C_{i}$ and $U_{i}$. For each $\{i,j\}\in\binom{[p]}{2}$, let $Y_{\{i,j\}}\subseteq V(G)$ and $X_{\{i,j\}}\subseteq V(G)$ be the sets obtained by applying Lemma 7 to $C_{i}$, $U_{i}$, $Y_{i}$, and $C_{j}$, $U_{j}$, $Y_{j}$.
If any of these applications of Lemma 6 or Lemma 7 produces a $d$-packing of $k$ cycles in $G$, then $G$ satisfies (a). Otherwise, the unicycles $U_{1},\ldots,U_{p}$ and the sets
@@
satisfy the conditions of Lemma 5. In particular, $|X|<k\cdot(2k+s(k))+\binom{k}{2}\cdot s(k)$, as desired. ∎
We are now ready to prove (the following quantitative version of) Theorem 1.
@@
Let $k\geq 1$ and $d\geq 1$ be integers and let $G$ be a graph. If $G$ contains a $d$-packing of $k$ cycles then there is nothing to prove. Thus, assume the opposite. Let $\{\leavevmode\color[rgb]{0.76,0.13,0.28}C_{1},\ldots,C_{p}\}$ be a maximal $2d$-packing of cycles that are $d$-unicyclic in $G$. Since $G$ has no $d$-packing of $k$ cycles, $p<k$.
Consider the following iterative process, that constructs a sequence of cycles $\mathcal{D}^{\prime}:=\{D^{\prime}_{1},D^{\prime}_{2},\ldots\}$ such that $\mathcal{D}^{\prime}$ is a $d$-packing of cycles each having length at most $26d+$ 2. Let $q$ be a nonnegative integer and suppose we have already constructed $D^{\prime}_{1},\ldots,D^{\prime}_{q}$.
If every cycle in $G$ contains a vertex in $B_{G}(\bigcup_{i\in[p]}V(C_{i}),4d)\cup B_{G}(\bigcup_{i\in[q]}V(D^{\prime}_{i}),4d)$ then the process ends. Otherwise, let $D$ be a cycle with no vertex in $B_{G}(\bigcup_{i\in[p]}V(C_{i}),4d)\cup B_{G}(\bigcup_{i\in[q]}V(D^{\prime}_{i}),4d)$. Apply Lemma 4 to $D$ to find a cycle $D^{\prime}$ such that either (a) $V(D^{\prime})\subseteq B_{G}(V(D),2d)$and $D^{\prime}$ is $d$-unicyclic in $G$; or (b) $V(D^{\prime})\subseteq B_{G}(V(D),3d)$and $D^{\prime}$ is of length at most $26d+$ 2. In case Item a, $V(D)\cap B_{G}(\bigcup_{i\in[p]}V(C_{i}),4d)=\emptyset$ and $V(D^{\prime})\subseteq B_{G}(V(D),2d)$, which implies that $V(D^{\prime})\cap B_{G}(\bigcup_{i\in[p]}V(C_{i}),2d)=\emptyset$. Therefore, $\{C_{1},\ldots,C_{p},D^{\prime}\}$ is a $2d$-packing of $d$-unicyclic cycles in $G$, which contradicts the choice of $\{C_{1},\ldots,C_{p}\}$.
Therefore case Item b applies. By the choice of $D$, $\operatorname{dist}_{G}(V(D),V(D_{i}^{\prime}))>4d$ for each $i\in[q]$. Since $V(D^{\prime})\subseteq B_{G}(V(D),3d)$, this implies that $\operatorname{dist}_{G}(V(D^{\prime}),V(D^{\prime}_{i}))>d$ for each $i\in[q]$. Set $D^{\prime}_{q+1}:=D^{\prime}$. Thus $\{D^{\prime}_{1},\ldots,D^{\prime}_{q+1}\}$ is a $d$-packing of cycles in $G$. This completes the description of the process. Since $G$ has no $d$-packing of $k$ cycles, this process produces a family $\mathcal{D^{\prime}}$ of size $q<k$. (Possibly $\mathcal{D^{\prime}}=\emptyset$ and $q=0$.)
@@
$\leavevmode\color[rgb]{0.76,0.13,0.28}Y_{0}:=\bigcup_{i\in[q]}B_{G}(V(D^{\prime}_{i}),4d).$
Let $\leavevmode\color[rgb]{0.76,0.13,0.28}X_{0}:=\{x_{1},\ldots,x_{q}\}$ where $x_{i}$ is an arbitrary vertex of $D^{\prime}_{i}$ for each $i\in[q]$. Since $D_{i}^{\prime}$ has length at most $26d+$ 2, we have $V(D_{i}^{\prime})\subseteq B_{G}(x_{i},3d+1)$ for each $i\in[q]$. Therefore,
@@
$\leavevmode\color[rgb]{0.76,0.13,0.28}r:=6d\enspace.$
Apply Lemma 5 to $\mathcal{C}:=\{C_{1},\ldots,C_{p}\}$ to obtain unicyclic graphs $U_{1},\ldots,U_{p}$, a set $Y_{1}\subseteq V(G)$ and a set $X_{1}\subseteq V(G)$ such that
@@
    for each $i\in[p]$, $U_{i}$ is a $C_{i}$-rooted BFS-spanning unicycle of $G[B_{G}(V(C_{i}),r)]$;
@@
For each $i\in[p]$, let $y_{i}$ be an arbitrary vertex of $C_{i}$ and define $\leavevmode\color[rgb]{0.76,0.13,0.28}Y_{2}:=\leavevmode\color[rgb]{0.76,0.13,0.28}X_{2}:=\{y_{1},\ldots,y_{p}\}$. Define
$\leavevmode\color[rgb]{0.76,0.13,0.28}\widehat{Y}$ $:=\textstyle Y_{0}\cup Y_{1}\cup Y_{2},\ \textrm{and}$
$\leavevmode\color[rgb]{0.76,0.13,0.28}\widehat{X}$ $:=\textstyle X_{0}\cup X_{1}\cup X_{2}.$
@@
$\leavevmode\color[rgb]{0.76,0.13,0.28}F_{0}$ $:=\textstyle{G-\left(\widehat{Y}\cup\bigcup_{i\in[p]}B_{G}(V(C_{i}),r-2d)\right)}\text{, and}$
$\leavevmode\color[rgb]{0.76,0.13,0.28}F^{-}_{0}$ $:=\textstyle{G-\left(\widehat{Y}\cup\bigcup_{i\in[p]}B_{G}(V(C_{i}),r-d)\right)}\enspace.$
@@
(Note that possibly $i=j$.) Each good tuple $t:=(e,i,e^{\prime},j)$ defines a walk $\leavevmode\color[rgb]{0.76,0.13,0.28}W_{t}:=P_{1}eP_{0}e^{\prime}P_{2}$ in $G$ where
@@
We call the path $P_{0}$ the forest leg of $W_{t}$, $P_{1}$ the first leg of $W_{t}$, and $P_{2}$ the second leg of $W_{t}$. The walk $eP_{0}e^{\prime}$ is called the extended forest leg of $W_{t}$. Note that the first and second legs of $W_{t}$ each have exactly $r-d$ edges, while the forest leg of $W_{t}$ could be arbitrarily long.
A tuple $t$ is admissible if $t$ is good and $B_{G}(V(W_{t}),d)\cap\widehat{Y}=\emptyset$. The following claim allows us to finish the proof by finding a set of balls that intersect the extended forest leg of each admissible tuple.
@@
We may assume that $C$ contains no vertex in $\widehat{Y}$ since, otherwise, there is nothing to prove. Since $F_{0}^{-}$ is an induced forest in $G$, $C$ must contain some vertex not in $F_{0}^{-}$. Therefore $C$ contains a vertex in $B_{G}(V(C_{i}),r-d)$ for some $i\in\{1,\ldots,p\}$. By construction, $G[B_{G}(V(C_{i}),r-d)]-Y_{1}$ is unicyclic and, if it contains any cycle then it contains $C_{i}$. Therefore $G[B_{G}(V(C_{i}),r-d)]-(Y_{1}\cup\{y_{i}\})$ is a forest. Since $C$ contains no vertex in $Y_{1}\cup\{y_{i}\}\subseteq\widehat{Y}$, $C$ must also contain a vertex not in $B_{G}(V(C_{i}),r-d)$.
Therefore, $C$ contains an edge $v_{0}v_{1}$ with $v_{0}\in B_{G}(V(C_{i}),r-d)$ and $v_{1}\not\in B_{G}(V(C_{i}),r-d)$. Since $d\geq 1$, $v_{1}\in B_{G}(V(C_{i}),r)$. Now, if $v_{1}\in B_{G}(V(C_{j}),r)$ for some $j\in[p]\setminus\{i\}$, then $v_{1}\in B_{G}(V(C_{j}),r)\cap B_{G}(V(C_{i}),r)\subseteq Y_{1}\subseteq\widehat{Y}$, a contradiction. Therefore, $v_{1}\in V(F_{0}^{-})$, so $v_{0}v_{1}$ is an $i$-exit edge.
Consider the maximal path $v_{1},\ldots,v_{h-1}$ in $C$ that contains $v_{1}$ and that is contained in $F_{0}^{-}$. Let $v_{h}$ be the neighbour of $v_{h-1}$ in $C-\{v_{h-2}\}$. (Possibly $v_{h}=v_{0}$.) Since $v_{h}\notin\widehat{Y}$ and $v_{h}$ is not a vertex of $F_{0}^{-}$, $v_{h}\in B_{G}(V(C_{j}),r-d)$ for some $j\in\{1,\ldots,p\}$. Then $v_{h-1}v_{h}$ is a $j$-exit edge and $v_{0},\ldots,v_{h}$ is the extended forest leg of the good tuple $t:=(v_{0}v_{1},i,v_{h-1}v_{h},j)$. If $t$ is admissible, then there is nothing more to prove. If $t$ is not admissible, then $B_{G}(V(W_{t}),d)$ contains a vertex of $\widehat{Y}$, say $m$. Let $P_{0}=v_{1},\ldots,v_{h-1}$ be the forest leg of $W_{t}$, let $P_{1}$ be the first leg of $W_{t}$ and let $P_{2}$ be the second leg of $W_{t}$. If $m\in B_{G}(V(P_{1}),d)$ then there is $u\in V(P_{1})$ such that $\operatorname{dist}_{G}(m,u)\leq d$ and so $\operatorname{dist}_{G}(m,v_{0})\leq\operatorname{dist}_{G}(m,u)+\operatorname{dist}_{G}(u,v_{0})\leq d+(r-d)$. If $m\in B_{G}(V(P_{2}),d)$ then $\operatorname{dist}_{G}(m,v_{h})\leq r$. Similarly, if $m\in B_{G}(V(P_{0}),d)$ then $\operatorname{dist}_{G}(m,V(P_{0}))\leq d$. In each case $B_{G}(m,r)$ contains a vertex in $V(C)\supseteq\{v_{0},\ldots,v_{h}\}$, a contradiction. ∎
Let $t=(e,i,e^{\prime},j)$ be an admissible tuple with $P_{1}eP_{0}e^{\prime}P_{2}:=W_{t}$. Define the $(p+1)$-tuple $\leavevmode\color[rgb]{0.76,0.13,0.28}\Psi(t):=(\Psi_{0}(t),\Psi_{1}(t),\ldots,\Psi_{p}(t))$ where
$\leavevmode\color[rgb]{0.76,0.13,0.28}\Psi_{\ell}(t)$ :=$:=\begin{cases}B_{G}(V(P_{0}),d)&\textrm{if $\ell=0$,}\\ B_{G}(V(P_{1}),d)&\textrm{if $\ell=i$ and $\ell\neq j$,}\\ B_{G}(V(P_{2}),d)&\textrm{if $\ell\neq i$ and $\ell=j$,}\\ B_{G}(V(P_{1}),d)\cup B_{G}(V(P_{2}),d)&\textrm{if $\ell=i$ and $\ell=j$,}\\ \emptyset&\textrm{if $\ell\notin\{0,i,j\}$.}\end{cases}$
@@
$\leavevmode\color[rgb]{0.76,0.13,0.28}F_{i}$ $:=G[B_{G}(V(C_{i}),r)]-\widehat{Y},$
$\leavevmode\color[rgb]{0.76,0.13,0.28}F_{i}^{-}$ $:=G[B_{G}(V(C_{i}),r-d)]-\widehat{Y}.$

### v3

@@ 4 The Proof
The only remaining possibility is that $v^{\prime}$ is an internal vertex of $P_{x^{\prime}}$ for some $x^{\prime}\in I$. Since every internal vertex of $P_{x^{\prime}}$ has degree $2$ in $G^{\prime}$, this implies that $P_{x^{\prime}}\subseteq D^{\prime}$. Let $w^{\prime}$ be the endpoint of $P_{x^{\prime}}$ in $C_{1}$. Since $^{\prime}\in I\subseteq V(H), xx^{\prime}$ is interesting. Therefore $P_{x^{\prime}}\subseteq P_{1,x^{\prime}}\cup P_{2,x^{\prime}}$ contains no vertex in $Y_{1}\cup Y_{2}$.
@@
First suppose that $v^{\prime}\in V(Q^{\prime})$. Let $Q$ be the subpath of $Q^{\prime}$ from $w^{\prime}$ to $v^{\prime}$. Since $Q\subseteq Q^{\prime}\subseteq P_{x^{\prime}}$ contains no vertex in $Y_{1}$ and $V(Q^{\prime})\subseteq B_{G}(V(C_{1}),r)$, we have that $Q\subseteq U_{1}$. Since $U_{1}$ is a $C_{1}$-rooted spanning BFS-unicycle of $G_{1,r}$, $\operatorname{dist}_{G}(w^{\prime},v^{\prime})=\operatorname{dist}_{U_{1}}(w^{\prime},v^{\prime})=\operatorname{len}(Q)$. Observe that $Q$ is a shortest path in $G$ between $V(C_{1})$ and $v^{\prime}$, and thus $\operatorname{len}(Q)\leq\operatorname{len}(P)\leq d$. It follows that $Q\subseteq G[B_{G}(V(C_{1}),d)]$. Since $C_{1}$ is $d$-unicyclic, all paths from $V(C_{1})$ to $v^{\prime}$ in $G[B_{G}(V(C_{1}),d)]$ contain $w^{\prime}$. In particular, $P\subseteq G[B_{G}(V(C_{1}),d)]$ is a path from $v\in V(C_{1})$ to $v^{\prime}$, so $w^{\prime}\in V(P)$. Since $V(P)\cap V(C_{1})=\{v\}$, this implies that $v=w^{\prime}$. However $v=w^{\prime}\in V(P_{x^{\prime}})\subseteq V(D^{\prime})$. This implies that $v$ is a vertex of both $D$ and $D^{\prime}$, which contradicts the fact that $D$ and $D^{\prime}$ are vertex-disjoint.
@@
$\geq 2(r+1)-2d+(2d+1)\geq 2r+3.$
(The second inequality follows from $\operatorname{dist}_{G}(V(C_{1}),q^{\prime})\leq\operatorname{dist}_{G}(V(C_{1}),v^{\prime})+\operatorname{dist}_{G}(v^{\prime},q^{\prime})$ and $\operatorname{dist}_{G}(V(C_{1}),V(C_{2}))\leq\operatorname{dist}_{G}(V(C_{1}),v^{\prime})+\operatorname{dist}_{G}(v^{\prime},V(C_{2}))$.) This contradicts the fact that $\operatorname{len}(P_{x^{\prime}})\leq\operatorname{len}(P_{1,x^{\prime}})+\operatorname{len}(P_{2,x^{\prime}})\leq 2r$. This completes the proof (assuming $|I|\geq\tfrac{1}{2}s(k)$) that outcome ˜a holds.
@@
$X:=\bigcup_{y\in I}\left(V(P_{y})\cap(V(C_{1})\cup V(C_{2}))\right),$
be the set of degree-$3$ vertices in $G^{\prime}$. Note that $|X|=2\cdot|I|<s(k)$. We show that $X$ and $Y$ satisfy ˜b.
Consider some vertex $y\in Y$. If $y$ is a $U_{i}$-descendant of some vertex in $Y_{i}$ for some $i\in[2]$, then $y$ satisfies Item˜x. Therefore, we assume that this is not the case, which implies that $y$ is interesting, so $y\in V(H)$. Since $I$ is a maximal independent set in $H$, the set $I$ is also a dominating set. Therefore, either $y\in I$ or $y$ has a neighbor in $I$. In $G$, this translates into the statement: there exists $y^{\prime}\in I$ such that $\operatorname{dist}_{G}(V(P_{y}),V(P_{y^{\prime}}))\leq d$. (This includes the case $y=y^{\prime}$, which applies when $y\in I$.)
@@
We have just shown that $X$ and $Y$ satisfy ˜ii. Furthermore, $Y=V(G_{1,r})\cap V(G_{2,r})$. Thus, the sets $X$ and $Y$ satisfy the conditions of outcome ˜b. ∎
Proof of Lemma˜5.
Let $\mathcal{C}$ be a $2d$-packing of $d$-unicyclic cycles in $G$. If $|\mathcal{C}|\geq k$, then ˜a holds. Thus, let $\mathcal{C}:=\{C_{1},\ldots,C_{p}\}$ and we assume that $p<k$.
For each $i\in[p]$, let $U_{i}$ be a spanning BFS-unicycle of $G[B_{G}(V(C_{i}),r)]$. For each $i\in[p]$, we apply Lemma˜6 to $C_{i}$ and $U_{i}$. If any of these applications of Lemma˜6 produces a $d$-packing of $k$ cycles in $G$, then $G$ satisfies (a) and there is nothing to prove. Thus, for each $i\in[p]$ we obtain $Y_{i}\subseteq V(G)$ and $X_{i}\subseteq V(G)$ as in the statement of Lemma˜6. For each $\{i,j\}\in\binom{[p]}{2}$, we apply Lemma˜7 to $C_{i}$, $U_{i}$, $Y_{i}$, and $C_{j}$, $U_{j}$, $Y_{j}$. If any of these applications of Lemma˜7 produces a $d$-packing of $k$ cycles in $G$, then $G$ satisfies (a) and there is nothing to prove. Thus, for each $\{i,j\}\in\binom{[p]}{2}$ we obtain $Y_{\{i,j\}}\subseteq V(G)$ and $X_{\{i,j\}}\subseteq V(G)$ as in the statement of Lemma˜7. Now, the unicycles $U_{1},\ldots,U_{p}$ and the sets
@@
satisfy the conditions of Lemma˜5. In particular, $|X|<k\cdot(2k+s(k))+\binom{k}{2}\cdot s(k)$, as desired. ∎
We are now ready to prove (the following quantitative version of) Theorem˜1.
@@
Let $k\geq 1$ and $d\geq 1$ be integers and let $G$ be a graph. If $G$ contains a $d$-packing of $k$ cycles then there is nothing to prove. Thus, assume the opposite. Let $\{\color[rgb]{0.76,0.13,0.28}C_{1},\ldots,C_{p}\}$ be a maximal $2d$-packing of cycles that are $d$-unicyclic in $G$. Since $G$ has no $d$-packing of $k$ cycles, $p<k$.
Consider the following iterative process, that constructs a sequence of cycles $\mathcal{D}^{\prime}:=\{D^{\prime}_{1},D^{\prime}_{2},\ldots\}$ such that $\mathcal{D}^{\prime}$ is a $d$-packing of cycles each having length at most $6d+2$. Let $q$ be a nonnegative integer and suppose we have already constructed $D^{\prime}_{1},\ldots,D^{\prime}_{q}$.
If every cycle in $G$ contains a vertex in $B_{G}(\bigcup_{i\in[p]}V(C_{i}),4d)\cup B_{G}(\bigcup_{i\in[q]}V(D^{\prime}_{i}),4d)$ then we let the process end. Otherwise, let $D$ be a cycle with no vertex in $B_{G}(\bigcup_{i\in[p]}V(C_{i}),4d)\cup B_{G}(\bigcup_{i\in[q]}V(D^{\prime}_{i}),4d)$. Apply Lemma˜4 to $D$ to find a cycle $D^{\prime}$ such that either (a) $V(D^{\prime})\subseteq B_{G}(V(D),2d)$and $D^{\prime}$ is $d$-unicyclic in $G$; or (b) $V(D^{\prime})\subseteq B_{G}(V(D),3d)$and $D^{\prime}$ is of length at most $6d+2$. In case Item˜a, $V(D)\cap B_{G}(\bigcup_{i\in[p]}V(C_{i}),4d)=\emptyset$ and $V(D^{\prime})\subseteq B_{G}(V(D),2d)$, which implies that $V(D^{\prime})\cap B_{G}(\bigcup_{i\in[p]}V(C_{i}),2d)=\emptyset$. Therefore, $\{C_{1},\ldots,C_{p},D^{\prime}\}$ is a $2d$-packing of $d$-unicyclic cycles in $G$, which contradicts the choice of $\{C_{1},\ldots,C_{p}\}$.
Therefore case Item˜b applies. By the choice of $D$, $\operatorname{dist}_{G}(V(D),V(D_{i}^{\prime}))>4d$ for each $i\in[q]$. Since $V(D^{\prime})\subseteq B_{G}(V(D),3d)$, this implies that $\operatorname{dist}_{G}(V(D^{\prime}),V(D^{\prime}_{i}))>d$ for each $i\in[q]$. Set $D^{\prime}_{q+1}:=D^{\prime}$. Thus $\{D^{\prime}_{1},\ldots,D^{\prime}_{q+1}\}$ is a $d$-packing of cycles in $G$. This completes the description of the process. Since $G$ has no $d$-packing of $k$ cycles, this process produces a family $\mathcal{D^{\prime}}$ of size $q<k$. (Possibly $\mathcal{D^{\prime}}=\emptyset$ and $q=0$.)
@@
$\color[rgb]{0.76,0.13,0.28}Y_{0}:=\bigcup_{i\in[q]}B_{G}(V(D^{\prime}_{i}),4d).$
Let $\color[rgb]{0.76,0.13,0.28}X_{0}:=\{x_{1},\ldots,x_{q}\}$ where $x_{i}$ is an arbitrary vertex of $D^{\prime}_{i}$ for each $i\in[q]$. Since $D_{i}^{\prime}$ has length at most $6d+2$, we have $V(D_{i}^{\prime})\subseteq B_{G}(x_{i},3d+1)$ for each $i\in[q]$. Therefore,
@@
$\color[rgb]{0.76,0.13,0.28}r:=6d.$
Apply Lemma˜5 to $\mathcal{C}:=\{C_{1},\ldots,C_{p}\}$ to obtain unicyclic graphs $U_{1},\ldots,U_{p}$, a set $Y_{1}\subseteq V(G)$ and a set $X_{1}\subseteq V(G)$ such that
@@
    for each $i\in[p]$, $U_{i}$ is a $C_{i}$-rooted spanning BFS-unicycle of $G[B_{G}(V(C_{i}),r)]$;
@@
For each $i\in[p]$, let $y_{i}$ be an arbitrary vertex of $C_{i}$ and define $\color[rgb]{0.76,0.13,0.28}Y_{2}:=\color[rgb]{0.76,0.13,0.28}X_{2}:=\{y_{1},\ldots,y_{p}\}$. Define
$\color[rgb]{0.76,0.13,0.28}\widehat{Y}$ $:=\textstyle Y_{0}\cup Y_{1}\cup Y_{2},\ \textrm{and}$
$\color[rgb]{0.76,0.13,0.28}\widehat{X}$ $:=\textstyle X_{0}\cup X_{1}\cup X_{2}.$
@@
$\color[rgb]{0.76,0.13,0.28}F_{0}$ $:=\textstyle{G-\left(\widehat{Y}\cup\bigcup_{i\in[p]}B_{G}(V(C_{i}),r-2d)\right)}\text{, and}$
$\color[rgb]{0.76,0.13,0.28}F^{-}_{0}$ $:=\textstyle{G-\left(\widehat{Y}\cup\bigcup_{i\in[p]}B_{G}(V(C_{i}),r-d)\right)}.$
@@
(Note that possibly $i=j$.) Each good tuple $t:=(e,i,e^{\prime},j)$ defines a walk $\color[rgb]{0.76,0.13,0.28}W_{t}:=P_{1}eP_{0}e^{\prime}P_{2}$ in $G$ where
@@
Figure 3: A good tuple $t=(e,i,e^{\prime},j)$ and the walk $W_{t}$.
We call the path $P_{0}$ the forest leg of $W_{t}$, $P_{1}$ the first leg of $W_{t}$, and $P_{2}$ the second leg of $W_{t}$. The walk $eP_{0}e^{\prime}$ is called the extended forest leg of $W_{t}$. Note that the first and second legs of $W_{t}$ each have exactly rdr-d edges, while the forest leg of $W_{t}$ could be arbitrarily long.
A tuple $t$ is admissible if $t$ is good and $B_{G}(V(W_{t}),d)\cap\widehat{Y}$$\emptyset$. The following claim allows us to finish the proof by finding a set of balls that intersect the extended forest leg of each admissible tuple.
@@
We may assume that $C$ contains no vertex in $B_{G}(\widehat{Y},r)$ since, otherwise, there is nothing to prove. Since $F_{0}^{-}$ is an induced forest in $G$, $C$ must contain some vertex not in $F_{0}^{-}$. Therefore $C$ contains a vertex in $B_{G}(V(C_{i}),r-d)$ for some $i\in\{1,\ldots,p\}$. Recall that $Y_{1}$ was obtained from applying Lemma˜5 and that $G[B_{G}(V(C_{i}),r-d)]-Y_{1}$ is unicyclic by (b). Thus, if $G[B_{G}(V(C_{i}),r-d)]-Y_{1}$ contains any cycle then this cycle is $C_{i}$. Therefore $G[B_{G}(V(C_{i}),r-d)]-(Y_{1}\cup\{y_{i}\})$ is a forest. Since $C$ contains no vertex in $Y_{1}\cup\{y_{i}\}\subseteq\widehat{Y}$, $C$ must also contain a vertex not in $B_{G}(V(C_{i}),r-d)$.
Therefore, $C$ contains an edge $v_{0}v_{1}$ with $v_{0}\in B_{G}(V(C_{i}),r-d)$ and $v_{1}\not\in B_{G}(V(C_{i}),r-d)$. Since $d\geq 1$, $v_{1}\in B_{G}(V(C_{i}),r)$. Now, if $v_{1}\in B_{G}(V(C_{j}),r)$ for some $j\in[p]\setminus\{i\}$, then $v_{1}\in B_{G}(V(C_{j}),r)\cap B_{G}(V(C_{i}),r)\subseteq Y_{1}\subseteq\widehat{Y}$ by (3), a contradiction. Therefore, $v_{1}\in V(F_{0}^{-})$, so $v_{0}v_{1}$ is an $i$-exit edge.
Consider the maximal path $v_{1},\ldots,v_{h-1}$ in $C$ that contains $v_{1}$ and that is contained in $F_{0}^{-}$. Let $v_{h}$ be the neighbour of $v_{h-1}$ in $C-\{v_{h-2}\}$. (Possibly $v_{h}=v_{0}$.) Since $v_{h}\notin\widehat{Y}$ and $v_{h}$ is not a vertex of $F_{0}^{-}$, $v_{h}\in B_{G}(V(C_{j}),r-d)$ for some $j\in\{1,\ldots,p\}$. Then $v_{h-1}v_{h}$ is a $j$-exit edge. Since $F_{0}^{-}$ is a forest, $v_{1},\ldots,v_{h-1}$ is the unique path between $v_{1}$ and $v_{h-1}$ in $F_{0}^{-}$, and it follows that $v_{0},\ldots,v_{h}$ is the extended forest leg of $W_{t}$ for the good tuple $t:=(v_{0}v_{1},i,v_{h-1}v_{h},j)$. If $t$ is admissible, then there is nothing more to prove. If $t$ is not admissible, then $B_{G}(V(W_{t}),d)$ contains a vertex of $\widehat{Y}$, say $m$. Let $P_{0}=v_{1},\ldots,v_{h-1}$ be the forest leg of $W_{t}$, let $P_{1}$ be the first leg of $W_{t}$ and let $P_{2}$ be the second leg of $W_{t}$. If $m\in B_{G}(V(P_{1}),d)$ then there is $u\in V(P_{1})$ such that $\operatorname{dist}_{G}(m,u)\leq d$ and so $\operatorname{dist}_{G}(m,v_{0})\leq\operatorname{dist}_{G}(m,u)+\operatorname{dist}_{G}(u,v_{0})\leq d+(r-d)$. If $m\in B_{G}(V(P_{2}),d)$ then $\operatorname{dist}_{G}(m,v_{h})\leq r$. Similarly, if $m\in B_{G}(V(P_{0}),d)$ then $\operatorname{dist}_{G}(m,V(P_{0}))\leq d$. In each case $B_{G}(m,r)$ contains a vertex in $V(C)\supseteq\{v_{0},\ldots,v_{h}\}$, a contradiction. ∎
Let $t=(e,i,e^{\prime},j)$ be an admissible tuple with $P_{1}eP_{0}e^{\prime}P_{2}:=W_{t}$. Recall that $p=|\mathcal{C}|$. Define the $(p+1)$-tuple $\color[rgb]{0.76,0.13,0.28}\Psi(t):=(\Psi_{0}(t),\Psi_{1}(t),\ldots,\Psi_{p}(t))$ where
[…]
@@
$\color[rgb]{0.76,0.13,0.28}F_{i}$ $:=G[B_{G}(V(C_{i}),r)]-\widehat{Y}$,
$\color[rgb]{0.76,0.13,0.28}F_{i}^{-}$ $:=G[B_{G}(V(C_{i}),r-d)]-\widehat{Y}$.